Break-Even Sales and Cost-Volume-Profit Chart For the coming year, Sorkin Company anticipates a unit selling price of $98, a unit variable cost of $49, and fixed costs of $396,900 Required: 1. Compute the anticipated break-even sales in units units 2. Compute the sales (units) required to realize income from operations of $196,000. units 3. Construct a cost-volume-profit chart, assuming maximum sales of 16,200 units within the relevant range. From your chart, indicate whether each of the following sales levels would produce a profit, a loss, or break-even. $1,107,400 $989,800 $793,800 $597,800 $480,200 4. Determine the probable income (loss) from operations if sales total 13,000 units. If required, use the minus sign to indicate a loss.
Answer #1
Req 1:
Selling pirce: $ 98 per unit
Variable cost: $ 49 per unit
Fixed cost: $ 396900
contribution margin per unit = Selling price - variable cost
(98 - 49 ) = $ 49 per unit
Break even point in units = Fixed cost/ Contribution margin per unit
($396900 /$49) = 8,100 units
Req 2:Sales units to realize income of $ 196000
Desired profit: $196000
Desired contribution: $196000+ $396900 = $ 592900
Sales unit to achieve $196000 income: Desired contribution/ Contribution margin per unit
($592900/ 49) = 12100 units
Req 3:
Break even units: 8100 units
Break even in $ (8100 units @ 98)= $793800
This represent the sales at $793800 will be break even sales.
Sales below $793800, there will be a loss
Sales above $ 793800, there will be a profit for the business
Therefore,
Sales level Effects
$1,107,400 Profits
$989,800 Profits
$793,800 Break even
$597,800 Loss
$480,200 Loss
Req 4:
Sales units: 13,000 units
Contribution margin per unit: $ 49 per unit
Total contribution margin earned: 13,000 units @49 = $637000
Net incoem = Contribution earned- Fixed cost
(637000 -396900) = 240100
Profit $ 240100